# Inferring resource competition in microbial communities from time series

### v2

@@ II.1 Environment-mediated consumer-resource model
The consumer-resource model (1)-(2), environmental drive (5)-(7), and resource guild structure $r_{i\alpha}$ among consumers, completely define the model. To illustrate its behavior and gain intuition, we first simulated a small network with $N=6$ species and $M=30$ resources, with a guild-structure bias of $q=0.4$. The resource-utilization matrix $\mathbf{G}\mathbf{G}^{\intercal}$ is given by Fig. 2(a). The specific set of parameters is given in Methods IV.1. The first $N_{a}=3$ species form one guild, while the remaining $N_{b}=3$ species form the other. To study the impact of environmental drive on species abundances, we simulate the ECRM under sinusoidal and OU environmental drives each sampled at three different ranges of timescales – slow, fast, and with a mixture of timescales (Fig. 2(b)). Example traces of pairs of species within the same guild show complex coupled dynamics depending on the environmental timescales (Fig. 2(d)).
@@ II.4 Spectral methods work for general ECRM systems with larger sizes
By the choice of scaling for $p_{\rm in}$ and $p_{\rm out}$, the $\mathbf{G}\mathbf{G}^{\intercal}$ matrix is sparse for all system sizes. We can set a threshold on our measures to perform link detection, i.e. identify whether there exists an effective interaction between species due to competition for the same resource. As shown in Fig. 5 for $q=0.5$ (and SI Fig. S1 for $q=0.4$), for all the three system sizes tested, the total coherence remains the best among the four predictors. The summary statistics in Fig. S2 show that the total coherence always performs the best among the four predictors, although all get worse as the system size increases. Guild detection (Fig. S2(c)) has a better AUC than link detection (Fig. S2(b)), especially for large system sizes, thanks to error correction afforded by single-linkage clustering.
@@ II.6 Reconstructing guild structure from relative species abundances
II.7 Applying temporal metrics to marine data
@@
Comparing the pairwise metrics of the species abundances and the pairwise phylogenetic distance (Fig. 7(a)), all four pairwise metrics exhibit a weak negative correlation with the phylogenetic distance, indicating that more closely related taxa are more likely to share resource utilization capabilities, which is consistent with the literature sireci_environmental_2023 ; lemos_phylogeny_2024 . After binning the phylogenetic distances using regular intervals and computing the mean and average of the pairwise metrics in each bin, the trend in negative correlations becomes more visible (Fig. 7(b), colored bars). To examine whether this negative correlation is statistically significant, we created a null model by destroying any phylogenetic correlation by randomly shuffling the indices of the time series before computing the pairwise metrics. The randomization is performed $10^{5}$ times to collect statistics. As shown by the gray error bars in Fig. 7(b), the null model results in flat pairwise metrics for all metrics, as a function of the phylogenetic distance. At large phylogenetic distances, the data and the index-shuffled null model return the same mean and variance, as indicated by the overlaying errorbars in Fig. 7(b).
Among the four metrics, total coherence exhibits the biggest difference between the data and the null model for small phylogenetic distances, with a $p$-value of $5\times 10^{-5}$ for the test of the mean in the first bin against the index-shuffled null model (see Methods IV.4 for details). These results indicate that even for data collected experimentally in nature, total coherence works well in distinguishing structures in OTUs sharing common resources.
@@
As shown by Fig. 7(c), different metrics yield different predicted community structures. Clustering based on coherence gives two non-trivial clusters that exclusively contain OTUs of the same class, which does not occur for other metrics. The equal-time correlation $\mathbf{C}(0)$ and its absolute value $\lvert\mathbf{C}(0)\rvert$ return similar networks. Meanwhile, the total absolute CPSD gives a different network with one giant component and many singletons. This further suggests that the coherence metrics provides a useful predictor of species functional relationships.
Figure 7: Pairwise metrics of the relative abundances of marine eukaryotic OTUs correlate with phylogeny. (a) Scatterplot of pairwise metrics of relative abundance exhibit a weak anti-correlation with the pairwise sequence distance (Jukes-Cantor distance). Data are from marine_data (downsampled to $N=31$, see text). Pairwise measures include the equal-time correlation $\mathbf{C}(0)$ and its absolute value $\lvert\mathbf{C}(0)\rvert$, the total magnitude of CPSD, and total coherence. Spearman’s rank correlation coefficient $r$ and the $p$-value are given in each panel. (b) Binned average of the pairwise metrics vs. pairwise sequence distances. Binning is taken at regular intervals of the sequence distance. The mean across each bin is plotted with a cross symbol, and error bars given the standard deviation within each bin. In grey are shown the results from a randomized control, where the order of indices of eukaryotic OTUs are shuffled randomly (number of shuffling = $10^{5}$). The p-value for the mean of the pairwise metrics in the first bin to be different from the index shuffled null model is given in each panel (see Methods IV.4 for details). (c) Resource overlap networks inferred using the four pairwise metrics, with threshold set to produce $k=13$ clusters including singletons. The phylum and class of each OTU is given by respectively the color and the shape of the node, and in the legend with the format phylum;class.
@@ III Discussion
Within this dynamic resource-centric picture, our second key innovation was to move beyond simple correlations ($\mathbf{C}(0)$) by employing spectral methods that utilize full temporal information. Our method stands in contrast to covariance-based approaches spieceasi ; friedman_inferring_2012 that estimate covariance or inverse covariance matrices to infer associations but do not utilize dynamics. One exception is a recent study that employs temporal delays and Granger causality to infer associations mainali_detecting_2019 . Similarly, spectral metrics (CPSD and coherence) provide a high-fidelity picture of resource overlaps in communities by extracting information from multiple temporal delays. We expect these methods will prove powerful for inferring structure in complex communities where the acquisition of high-resolution time series has become commonplace.
@@
Using this approach, our analysis of eukaryotic microbial dynamics in a high-resolution marine dataset showed promising results in two domains. First, we observed that spectral methods outperform simple correlations in detecting statistically significant resource overlap between phylogenetically related taxa (Fig. 7B). Second, it is enticing that these methods yield distinct network architectures from the same data (Fig. 7C). An important avenue for future work is to more carefully vet the predicted associations from these metrics with synthetic communities.
@@ IV.1 Initialize parameters of ECRMs
For all system sizes, the parameters of the ECRM systems are set as the following. The uptake rate for non-zero resource intake is set to $r=0.1$. The yield $\gamma_{i\alpha}$ is given drawn from a Gaussian distribution with mean 1 and standard deviation 1/6. Redraws are performed to ensure all yields are non-negative. Death rate of the species is set to $d_{x}=0.3$, while depletion rate of the resource is set to $d_{R}=0.5$. For sinusoidal drives, the timescales of the environmental fluctuation, $\omega_{\alpha}$, are chosen randomly from a log-uniform distribution with a pre-determined range to represent the mixture of multiple timescales in real environment and to avoid resonance. The phase $\phi_{\alpha}$ is chosen from a uniform random distribution between 0 and $2\pi$. For OU environmental drives, thanks to its stochastic nature, we can choose the intrinsic frequencies $\omega_{\alpha}$ to be the same for all resources $\alpha$.
@@ IV.2 Finite size scaling in ECRMs
Case 1: species $i$ and $j$ belong to the same community. The probability of that there is a no link between two species is equal to the probability that there is no common resource between the two, which is
@@
Case 2: species $i$ and $j$ do not belong to the same community. The probability that there is no link is
@@
IV.3 Measure sequence distance in marine eukaryotic OTUs
@@
IV.4 Statistical tests for binned pairwise metrics against index shuffled null model
@@ Acknowledgments
Figure S2: Coherence and CPSD outperforms equal-time correlation measures in recovering the resource-utilization overlap matrix $\mathbf{G}\mathbf{G}^{\intercal}$ for all sizes of systems. (a) Spearman’s rank correlation coefficient between the four considered pairwise observables (equal time correlation $\mathbf{C}(0)$ and its absolute value $\lvert\mathbf{C}(0)\rvert$, the total magnitude of CPSD $\mathcal{P}$, and the total coherence $\mathcal{TC}$) and the elements of the $\mathbf{G}\mathbf{G}^{\intercal}$ matrix, for three families of ECRM systems with $N=6,24,96$ defined in Table 1. Error bars represent standard deviation across 100 random realizations of species abundances for systems with $N=6$ and with $N=24$ (10 random ECRM systems each with 10 random OU environmental drives with $\omega_{\alpha}=1$), and across 10 random realizations of species abundances for systems with $N=96$ (10 random ECRM systems each with 1 random OU environmental drive with $\omega_{\alpha}=1$). (b) Area under curve (AUC) for link detection for the same systems as in (a). Each AUC is computed for the same ECRM system, averaged over the ROC of all 10 random OU drives. (c) AUC for guild detection after single linkage clustering for the same systems as in (a).
@@
Figure S5: Pairwise measurements of the 93-day time series of relative abundance of eukaryotic OTUs in the Martin-Platero et al. coastal plankton dataset (Fig. 7, data from marine_data , downsampled to $N=31$). (a) Equal-time correlation $\mathbf{C}(0)$ and its absolute value $\lvert\mathbf{C}(0)\rvert$, the total magnitude of CPSD, and the total coherence of the time-series of relative abundance of marine eukaryotic OTUs ($N=31$). The phylum and the class of each OTU is indicated by an unique color and shape (given in legend at bottom). (b) The adjacency matrix formed by thresholding the pairwise measurements in panel a to form $k=13$ clusters (as in Fig. 7c and described in text). Black indicates a link, while white indicates no link.

### v3

@@ II.1 Environment-mediated consumer-resource model
The consumer-resource model (1)-(2), environmental drive (5)-(7), and resource guild structure $r_{i\alpha}$ among consumers, completely define the model. To illustrate its behavior and gain intuition, we first simulated a small network with $N=6$ species and $M=30$ resources, with a guild-structure bias of $q=0.4$. The resource-utilization matrix $\mathbf{G}\mathbf{G}^{\intercal}$ is given by Fig. 2(a). The specific set of parameters is given in Methods. The first $N_{a}=3$ species form one guild, while the remaining $N_{b}=3$ species form the other. To study the impact of environmental drive on species abundances, we simulate the ECRM under sinusoidal and OU environmental drives each sampled at three different ranges of timescales – slow, fast, and with a mixture of timescales (Fig. 2(b)). Example traces of pairs of species within the same guild show complex coupled dynamics depending on the environmental timescales (Fig. 2(d)).
@@ II.4 Spectral methods work for general ECRM systems with larger sizes
By the choice of scaling for $p_{\rm in}$ and $p_{\rm out}$, the $\mathbf{G}\mathbf{G}^{\intercal}$ matrix is sparse for all system sizes. We can set a threshold on our measures to perform link detection, i.e. identify whether there exists an effective interaction between species due to competition for the same resource. As shown in Fig. 5 for $q=0.5$ (and SI Fig. S1 for $q=0.4$), for all the three system sizes tested, the total coherence remains the best among the four predictors. The summary statistics in Fig. S2 show that the total coherence always performs the best among the four predictors, although all get worse as the system size increases. For $q=0.5$, guild detection (Fig. S2(c)) has a better AUC than link detection (Fig. S2(b)), especially for large system sizes, thanks to error correction afforded by single-linkage clustering. For $q=0.4$, however, guild detection performance is worse at larger system size, as single-linkage clustering over-corrects given the presence of non-zero resource competition across different guilds (Fig. S1(d,i)). Alternatively, one can perform single-linkage clustering but only for one iteration, such that immediately adjacent neighbors are now linked. While suggesting more sophisticated community-detection methods, this middle ground provides a better recovery of the guild structure compared to link detection and guild detection with full single-linkage clustering (Fig. S1(e,j)).
@@ II.6 Reconstructing guild structure from relative species abundances
II.7 Reconstruction of guild structure with heterogeneity in guild size and evenness
So far we have examined ECRMs where guild size is uniform across a community. Here, we test whether the dynamical observables still perform well when there is heterogeneity in guild size or unevenness in species abundances.
SI Fig. S5 shows that the ability to successfully detect links remains high for communities comprised of guilds of variable size. Coherence remains the best predictor among the four pairwise measures. Guild detection, however, becomes worse when the ECRM allows for the sharing of resources across different guilds ($q=0.4$). Detectability depends on guild size: while link detectability is unaffected by guild structure, smaller guilds are harder to detect after single linkage clustering.
Another source of heterogeneity is species evenness, i.e. certain species can be more abundant in the community. To allow variability in species abundances, especially across different guilds, we adjust the number of resources associated with each guild. SI Fig. S6 shows that the local guild detection is improved by species unevenness, especially if species in small guilds are present at high relative abundance. While local detectability isn’t monotonic with species evenness, higher relative abundance improves AUC. Thus, a small guild (e.g., two species) in a community with larger guilds may be misidentified—unless its members are highly abundant.
Figure 7: Pairwise metrics of the relative abundances of marine eukaryotic OTUs correlate with phylogeny. (a) Relative abundance of the marine eukaryotic OTUs (downsampled to $N=31$, see text) over 93 days. Each color represents one of the 31 OTUs. Data are from marine_data . (b) Scatterplot of pairwise metrics of relative abundance exhibit a weak anti-correlation with the pairwise sequence distance (Jukes-Cantor distance). Pairwise measures include the equal-time correlation $\mathbf{C}(0)$ and its absolute value $\lvert\mathbf{C}(0)\rvert$, the total magnitude of CPSD, and total coherence. Spearman’s rank correlation coefficient $r$ and the $p$-value are given in each panel. (c) Binned average of the pairwise metrics vs. pairwise sequence distances. Binning is taken at regular intervals of the sequence distance. The mean across each bin is plotted with a cross symbol (connected with lines to guide the eye), and error bars given the standard deviation within each bin. In grey are shown the results from a randomized control, where the order of indices of eukaryotic OTUs are shuffled randomly (number of shuffling = $10^{5}$). The p-value for the mean of the pairwise metrics in the first bin to be different from the index shuffled null model is given in each panel (see Methods IV.6 for details). (d) Resource overlap networks inferred using the four pairwise metrics, with threshold set to produce $k=13$ clusters including singletons. The phylum and class of each OTU is given by respectively the color and the shape of the node, and in the legend with the format phylum;class.
II.8 Applying temporal metrics to marine data
@@
Comparing the pairwise metrics of the species abundances and the pairwise phylogenetic distance (Fig. 7(a)), all four pairwise metrics exhibit a weak negative correlation with the phylogenetic distance, indicating that more closely related taxa are more likely to share resource utilization capabilities, which is consistent with the literature sireci_environmental_2023 ; lemos_phylogeny_2024 . After binning the phylogenetic distances using regular intervals and computing the mean and average of the pairwise metrics in each bin, the trend in negative correlations becomes more visible (Fig. 7(b), colored bars). To examine whether this negative correlation is statistically significant, we created a null model by destroying any phylogenetic correlation by randomly shuffling the indices of the OTUs before computing the pairwise metrics. The randomization is performed $10^{5}$ times to collect statistics. As shown by the gray error bars in Fig. 7(b), the null model results in flat pairwise metrics for all metrics, as a function of the phylogenetic distance. At large phylogenetic distances, the data and the index-shuffled null model return the same mean and variance, as indicated by the overlaying errorbars in Fig. 7(b).
Among the four metrics, total coherence exhibits the biggest difference between the data and the null model for small phylogenetic distances, with a $p$-value of $5\times 10^{-5}$ for the test of the mean in the first bin against the index-shuffled null model (see Methods IV.6 for details). These results indicate that even for data collected experimentally in nature, total coherence works well in distinguishing structures in OTUs sharing common resources.
@@
As shown by Fig. 7(c), different metrics yield different predicted guild structures. Clustering based on coherence gives two non-trivial clusters that exclusively contain OTUs of the same class, which does not occur for other metrics. The equal-time correlation $\mathbf{C}(0)$ and its absolute value $\lvert\mathbf{C}(0)\rvert$ return similar networks. Meanwhile, the total absolute CPSD gives a different network with one giant component and many singletons. This further suggests that the coherence metrics provide a useful predictor of species functional relationships.
@@ III Discussion
Within this dynamic resource-centric picture, our second key innovation was to move beyond simple correlations ($\mathbf{C}(0)$) by employing spectral methods that utilize full temporal information. Our method stands in contrast to covariance-based approaches spieceasi ; friedman_inferring_2012 that estimate covariance or inverse covariance matrices to infer associations but do not utilize dynamics. One exception is a recent study that employs temporal delays and Granger causality to infer associations mainali_detecting_2019 . Similarly, spectral metrics (CPSD and coherence) provide a high-fidelity picture of resource overlaps in communities by extracting information from multiple temporal delays. We expect these methods will prove powerful for inferring structure in complex community where the acquisition of high-resolution time series has become commonplace.
@@
Using this approach, our analysis of eukaryotic microbial dynamics in a high-resolution marine dataset showed promising results in two domains. First, we observed that spectral methods outperform simple correlations in detecting statistically significant resource overlap between phylogenetically related taxa (Fig. 7(b)). Second, it is enticing that these methods yield distinct network architectures from the same data (Fig. 7(c)). An important avenue for future work is to more carefully vet the predicted associations from these metrics with synthetic communities.
@@ IV.1 Initialize parameters of ECRMs
For all system sizes, the parameters of the ECRM systems are set as follows. The uptake rate for non-zero resource intake is set to $r=0.1$. The yield $\gamma_{i\alpha}$ is given, drawn from a Gaussian distribution with mean 1 and standard deviation 1/6. Redraws are performed to ensure all yields are non-negative. The death rate of the species is set to $d_{x}=0.3$, while the depletion rate of the resource is set to $d_{R}=0.5$. For sinusoidal drives, the timescales of the environmental fluctuation, $\omega_{\alpha}$, are chosen randomly from a log-uniform distribution with a pre-determined range to represent the mixture of multiple timescales in real environment and to avoid resonance. The phase $\phi_{\alpha}$ is chosen from a uniform random distribution between 0 and $2\pi$. For OU environmental drives, thanks to its stochastic nature, we can choose the intrinsic frequencies $\omega_{\alpha}$ to be the same for all resources $\alpha$.
@@ IV.2 Finite size scaling in ECRMs
Case 1: species $i$ and $j$ belong to the same guild. The probability that there is no link between two species is equal to the probability that there is no common resource between the two, which is
@@
Case 2: species $i$ and $j$ do not belong to the same guild. The probability that there is no link is
@@
IV.3 Initiating guild structure with guild-size variability
To introduce more realistic heterogeneity in the ECRMs, we introduce a tuning parameter $q_{N}$ which sets geometrically-spaced guild sizes, $N_{k}=q_{N}^{k-1}N_{1}$, and $\sum N_{k}=N$. Furthermore, to allow variability in species abundance, especially across different guilds, we introduce another tuning parameter $q_{M}$, which accordingly adjusts the number of resources associated within each guild, again by setting a geometric series, where $M_{k}=q_{M}^{k-1}M_{1}$, and $\sum M_{k}+N=M$. $N_{k}$’s and $M_{k}$’s are rounded up to the nearest integer. If the sum of the rounded up number is smaller than the total number of species $N$, the missing species is added to the largest guild. If the sum of the rounded up number is greater than the total number of species $N$, the additional species is subtracted from the smallest guild, unless if the smallest guild has only 2 species, then the additional species is subtracted from the second smallest guild, such that the minimum size of each guild is 2.
To give intuition, if $q_{N}=1$, all guild sizes are equal. The further $q_{N}$ is from 1, the greater the differences in guild sizes. If $q_{M}=q_{N}$, then on average each species is exposed to the same number of resources, and the species are even.
IV.4 Local measure of detectability
We introduce a local measure of detectability. Namely, for guild $k$, the true positive rate is the ratio between the number of detected adjacency links between all species present in guild $k$ and the number of possible pairs, $N_{k}(N_{k}-1)/2$. The guild-specific false positive rate is defined as the probability to falsely classify a link between a species $i$ in guild $k$ with a species $j$ outside of guild $k$.
IV.5 Measure sequence distance in marine eukaryotic OTUs
@@
IV.6 Statistical tests for binned pairwise metrics against index shuffled null model
@@ Acknowledgments
Figure S2: Coherence and CPSD outperforms equal-time correlation measures in recovering the resource-utilization overlap matrix $\mathbf{G}\mathbf{G}^{\intercal}$ for all sizes of systems. (a) Spearman’s rank correlation coefficient between the four considered pairwise observables (equal time correlation $\mathbf{C}(0)$ and its absolute value $\lvert\mathbf{C}(0)\rvert$, the total magnitude of CPSD $\mathcal{P}$, and the total coherence $\mathcal{TC}$) and the elements of the $\mathbf{G}\mathbf{G}^{\intercal}$ matrix, for three families of ECRM systems with $N=6,24,96$ defined in Table 1. Error bars represent standard deviation across 100 random realizations of species abundances for systems with $N=6$ and with $N=24$ (10 random ECRM systems each with 10 random OU environmental drives with $\omega_{\alpha}=1$), and across 10 random realizations of species abundances for systems with $N=96$ (10 random ECRM systems each with 1 random OU environmental drive with $\omega_{\alpha}=1$). (b) Area under curve (AUC) for link detection for the same systems as in (a). Each AUC is computed for the same ECRM system, averaged over the ROC of all 10 random OU drives. (c) AUC for guild detection after single linkage clustering for the same systems as in (a). Guild-structure bias is $q=0.5$.
@@
Figure S5: Link detection and guild detection for heterogeneous ECRM ecosystems with different guild sizes. The guild sizes follow a geometrical series with the common ratio set by the guild-size variability $q_{N}$ (see Methods). The average number of resources per species is tuned by $q_{S}=q_{M}/q_{N}$, where $q_{M}$ is the common ratio of the geometrical series of the number of resources associated with each guild. For this plot, $q_{S}=1$, which leads to each species being equally presented. Guild-structure bias is set to $q=0.4$. (a) Schematics for an ECRM with $N=24$ species, $M=120$ resources, and $k=4$ guilds. The guild structures at $q_{N}=1.1^{4}$. The number of species in each guild is $N_{1}=2$, $N_{2}=5$, $N_{3}=7$, $N_{4}=10$, respectively. (b) Averaged relative abundance for each species. Color represents distinct guilds. Error bars represent standard deviation across 10 random realizations of the $\mathbf{G}\mathbf{G}^{\intercal}$ matrix, each with 10 random OU environmental drives with $\omega_{\alpha}=1$. (c-d) ROC curve of link detection (c) and guild detection (d) for the four pairwise observables. Legend for the colors is given in panel (d). The black curve is the ROC curve if one uses the ground truth $\mathbf{G}\mathbf{G}^{\intercal}$ matrix as the binary classifier. Coherence remains the best predictor of the link and guild structure. (e-f) Guild-specific local detectability for link detection (e) and guild detection after single linkage clustering (f), using coherence as the binary classifier. Local detectability measures for each guild, the probability of two species correctly being identified in the guild (see detailed definition in Methods). Colors correspond to the four guilds with different sizes as given in panel (a). Although link detection does not depend on guild structure, guild detectability after single linkage clustering shows a considerable difference between guilds of different sizes. In particular, the smallest guild with $N_{1}=2$ becomes difficult to detect. (g-h) Link detection (g) and guild detection (h) for ECRM ecosystems with a range of guild-size variability. Here, we specify $q_{N}=1,1.1^{1},\dots,1.1^{4}$. $q_{S}=1$ always, such that the relative abundance for each species is even. Guild sizes are equal at $q_{N}=1$, which corresponds to the discussions in the main text.
Figure S6: Guild detection for heterogeneous ECRM ecosystems with different guild sizes and different species abundance. Detection of small guilds is rescued by large relative abundance in those guilds. Shown here are the results for an ECRM ecosystem with $N=24$ species, $M=120$ resources, $k=4$ guilds, and guild-size variability $q_{N}=1.1^{4}$ as exemplified by the schematics in Fig. S5(a). Averaged number of resources per species is tuned by $q_{S}=q_{M}/q_{N}$. (a) Relative abundance of the species in an example ECRM ecosystem, with $q_{M}=1.1^{-4}$, i.e. $q_{S}=q_{M}/q_{N}=1.1^{-8}$. This community is uneven. Species belonging to the smallest guild with size $N_{1}=1$ have the largest relative abundance. Colors represent distinct guilds. Error bars represent standard deviation across 10 random realization of the $\mathbf{G}\mathbf{G}^{\intercal}$ matrix, each with 10 random OU environmental drives with $\omega_{\alpha}=1$. (b) Guild-specific local detectability for guild detection, using coherence as the binary classifier. Curves with different colors correspond to the four different guilds. The detectability of the smallest guild with size $N_{1}=2$ (cyan) improves compared to the ECRM with the same guild sizes but even species abundance (Fig. S5)(f). (c) AUC for local guild detectability using coherence as the binary classifier (Area under the curves in panel (b)), for various species unevenness with the tuning parameter $q_{S}=1.1^{-12},1.1^{-11},\dots,1.1^{-1},1$. The value $q_{S}=1$ corresponds to results presented in Fig. S5, where species are even. (d) AUC for local guild detectability against the mean relative abundance of all species in specific guilds. As the relative abundance increases, the AUC also increases, suggesting that guilds are more detectable if species in the guild are more abundant. Specifically, at small values of $q_{S}$, i.e when the smaller guilds have large relative abundance and the large guilds have small relative abundance, it becomes more difficult to detect the larger guilds ($N_{4}=10$ curve in panel (c)), although the AUC is still greater than 0.5. Results in (c) and (d) are averages over 10 random realizations of the ECRM systems.
@@
Figure S7: Pairwise measurements of the 93-day time series of relative abundance of eukaryotic OTUs in the Martin-Platero et al. coastal plankton dataset (Fig. 7, data from marine_data , downsampled to $N=31$). (a) Equal-time correlation $\mathbf{C}(0)$ and its absolute value $\lvert\mathbf{C}(0)\rvert$, the total magnitude of CPSD, and the total coherence of the time-series of relative abundance of marine eukaryotic OTUs ($N=31$). The phylum and the class of each OTU is indicated by an unique color and shape (given in legend at bottom). (b) The adjacency matrix formed by thresholding the pairwise measurements in panel a to form $k=13$ clusters (as in Fig. 7(c) and described in text). Black indicates a link, while white indicates no link.